OpenStudy (anonymous):

Find dy/dx: y=e^(5-3lnx)

OpenStudy (anonymous):

I want to know how to differentiate 5-3lnx

OpenStudy (zepp):

For any dy/dx ln(x) = 1/x * f'(x)

OpenStudy (anonymous):

differentiation of 5 will be 0 and -3lnx will be -3/x

OpenStudy (zepp):

Ah, thought of a function, haha

OpenStudy (zepp):

\[\large \frac{d}{dx}\ln(x)=\frac{1}{x}\]

OpenStudy (anonymous):

yeah right:)

OpenStudy (zepp):

\[\large \frac{d}{dx}(5-3\ln x)=\frac{d}{dx}5-\frac{d}{dx}3\ln x=0-\frac{3}{x}=-\frac{3}{x}\]

OpenStudy (anonymous):

\[\large \frac{d}{dx}(e)^{f(x)} = (e)^{f(x)} \times \frac{d}{dx}(f(x))\]

OpenStudy (anonymous):

[e ^{5-3lnx}*(-3/x)\]

OpenStudy (anonymous):

Yes you are right..

OpenStudy (zepp):

Correct

OpenStudy (anonymous):

Get it thank you so much, i was just confused about what rule to use.

OpenStudy (anonymous):

Give the medal to zepp.. @jainilk

OpenStudy (zepp):

Apply the chain rule for this one

OpenStudy (anonymous):

I thought jainik asked the question.. Ha ha ha..